# When Tokens Talk Too Much: A Survey of Multimodal Long-Context Token Compression across Images, Videos, and Audios

### v3

@@ 4.4.1 Token Distillation
Token Turing Machines (TTMs) [104] maintain a compact external memory of summary tokens, sequentially compressing both new input tokens and memory at each timestep via a Transformer-based read/write mechanism, allowing scalable and efficient processing of long video sequences. BLIP-3-Video [105] introduces an explicit temporal encoder that abstracts hundreds of frame-level visual tokens into as few as 16–32 spatiotemporal tokens using learnable pooling and sequential models, enabling efficient video understanding with limited token usage. LinVT [108] proposes a plug-and-play Linear Video Tokenizer, which linearly aggregates frame-level visual tokens into a compact set of video tokens through spatio-temporal scoring, multi-scale pooling, and text-conditioned aggregation, enabling existing image-LLMs to efficiently process videos and dynamically extract question-relevant information. Long-VMNet [106] accelerates long-form video understanding by using a neural sampler to select discriminative visual tokens from clips and storing them in a fixed-size memory bank for each video; downstream queries are answered by processing only these memory tokens, greatly reducing computational cost while preserving key spatiotemporal information. STORM [107] inserts a Mamba-based [167] temporal encoder between the image encoder and LLM, using spatiotemporal scanning and pooling to inject temporal context into frame tokens and then aggressively compresses tokens by temporal and spatial pooling, enabling efficient long video understanding with minimal token loss. To understand more methods and applications of token distillation in video LLMs, please also refer to Section 3.4 for a detailed explanation.
4.4.2 Cross-Modal Selection
In video large language models (video LLMs), a query is commonly used to guide the selection of salient frames. In extreme cases, only a handful of frames are relevant to the posed question, allowing the tokens from the vast majority of remaining frames to be discarded. When dealing with an immense number of frames, finding query-relevant information can be akin to searching for a "needle in a haystack" for the LLM. Query-based token compression methods can pre-filter these query-relevant tokens, significantly alleviating the computational burden on the LLM.
@@ 4.4.2 Cross-Modal Selection
5 Audio-centric Token Compression
@@
The second category inherently reduces the number of tokens by optimizing the tokenizer structure and the design of the codebook. Nevertheless, detailed exploration of these specific design considerations falls outside the purview of this survey.
@@
5.1 Transformation-based audio-centric Compression
@@
5.1.1 Token Stacking
Similar to the pixel unshuffle operation in 2D image processing, this approach for audio LLMs token compression involves stacking multiple consecutive tokens along the hidden dimension of the token. This effectively reduces the total number of tokens. Notably, HTS-AT [109], an early example of audio token stacking for classification tasks within audio transformers, utilized 2D pixel-unshuffling on the 2D features extracted from Mel spectrograms to reduce audio tokens. More recent methods such as SLAM-ASR [110], LLaMA-Omni [111], Llama-AVSR [115] and others [172] have adopted this technique. Since these token stacking operations alter the hidden dimension, a MLP is typically used to realign the dimension for compatibility with other modalities.
5.1.2 Pooling
@@ 5.1.3 Temporal Convolution
5.2 Similarity-based audio-centric Compression
Similarity-based compression methods aim for each audio token to carry unique information rather than being overly redundant. Similarly to the ToMe [21] method used in vision transformers (ViT), A-ToMe [119] inserts a token merge module between the multihead self-attention (MHSA) and feed-forward network (FFN) in each layer. This module merges adjacent audio tokens that have high cosine similarity.
5.3 Attention-based audio-centric Compression
@@
5.3.1 Attention in Encoder
Top-K [120] operates within an audio spectrogram transformer block, directly retain only the top K audio tokens based on their attention scores. This prunes less attentive tokens, focusing on those with higher relevance as determined by the self-attention mechanism within the transformer.
5.3.2 Attention in Decoder
@@
5.4 Query-based audio-centric Compression
@@ 5.4.1 Token Distillation
Video-LLaMA [7] and SALMONN series [122, 123] employs an audio Q-former to transform variable-length audio inputs into a fixed-length sequence of learnable queries, thereby condensing audio information for the LLM. MMCE-Qformer [125] compresses acoustic information by utilizing learnable queries to extract global acoustic context from contextual audio embeddings. Concurrently, a cross-attention mechanism, guided by input text embeddings, captures local acoustic context relevant to each text token. This dual approach distills both broad and specific audio features into compact, text-relevant representations. MMS-LLaVA [126] reduces multimodal token length for efficient speech LLMs. It first halves the sequence length with an Early AV-Fusion Module, which combines visual and audio features. Subsequently, an AV Q-Former further compresses these fused features into a fixed number of queries, effectively capturing full speech context to bridge the token gap with text.
5.4.2 Cross-Modal Selection
@@
6 Discussions
6.1 Synergies and Distinctions with Other Compression Methods
@@
6.1.1 Weight-Focused Compression Methods
These methods mainly target the model dimension () by reducing the effective size or complexity of the model weights. Model Quantization reduces weight precision, directly impacting the memory associated with . A key limitation is that highly aggressive quantization (e.g., 4-bit) often compromises accuracy, meaning there’s no "free lunch" when it comes to achieving lossless performance. Furthermore, effectively accelerating these lower bit-rates often necessitates specialized hardware. Network Pruning removes redundant connections, effectively reducing the active parameters contributing to . For LLMs, aggressive structured pruning (e.g., beyond for downstream tasks) often leads to significant performance degradation or near-collapse due to the difficulty in preserving architectural integrity. Knowledge Distillation trains a smaller student model (with a smaller ) to mimic a larger teacher [183]. Its main limitation is the "knowledge gap", as the student may struggle to fully capture the teacher’s comprehensive knowledge, leading to performance disparities, especially on complex or out-of-distribution data. Low-Rank Factorization decomposes weight matrices into lower-rank approximations, thus reducing parameters related to . The challenge lies in finding an optimal low-rank approximation for diverse tasks without performance loss, as this is often task-dependent and complex to apply consistently across deep networks.
6.1.2 Token Compression
@@
6.1.3 Complementary Nature and Synergistic Gains
@@
6.2 Token Compression: Efficiency and Beyond
@@
6.2.1 Enhanced Modality Alignment
@@
6.2.2 Improved Information Representation
@@
Addressing these issues, the method classifications discussed earlier directly correspond to these types of data redundancy. Specifically, the transformation-based methods elaborated in Section 3.1, 4.1, 5.1, along with similarity-based approaches in Section 3.2, 4.2, 5.2, are effective in mitigating the feature redundancy. Furthermore, attention-based methods, as presented in Section 3.3, 4.3, 5.3, play a crucial role in minimizing attention computation redundancy. Lastly, query-based methods, detailed in Section 3.4, 4.4, 5.4, are designed to reduce task-irrelevant redundancy.
6.2.3 Enable One-Shot Long-Context Understanding
@@
6.3 Combining Different Token Compression Methods
In Section 6.2.2, we explored three distinct types of redundancy and the corresponding methods to reduce them. Since prune different forms of redundancy and can be applied at various stages, they seem orthogonal in their function. This raises a natural question: can we combine multiple token compression methods to achieve a synergistic effect?
Surprisingly, the answer is often no. Research has shown that a simple combination can lead to a phenomenon where the total effect is less than the sum of its parts. For example, the MoB [196] paper analyzed this in the context of visual token pruning, finding that the combining methods do not exhibit superiority. This is probably attributed to the difference in each compression technique. When methods target different redundancies, their compression metrics can be inconsistent, leading to conflicting pruning decisions and a suboptimal outcome. Therefore, while each method excels at its specific task, their combined use does not necessarily guarantee a better result and may even degrade performance.
6.4 Current Challenges
6.4.1 Performance Degradation
While token compression can effectively condense and abstract multimodal features, it also introduces a risk of performance degradation. Current research on visual MLLMs, for example, has shown that for models like LLaVA-OV-7B [2], near-lossless performance can be achieved by retaining as few as of the original tokens. However, performance declines sharply when the compression rate is pushed further. This challenge is even more pronounced for larger and more recent models such as Qwen2.5-VL [33], LLaVA-Video-7B [34] and LLaVA-OV-72B [2], where achieving lossless compression seems to be more difficult.
This increased difficulty may stem from the models’ enhanced representational capabilities. It has been suggested that less capable models are inherently less sensitive to information loss from aggressive compression, as their weaker understanding already struggles to fully process the complex, uncompressed data. In contrast, more sophisticated models, which possess a more nuanced and holistic comprehension of multimodal tokens, are more susceptible to the subtle degradation caused by compression. For these models, achieving high performance requires a far more delicate and precise approach to preserve the token.
6.4.2 Task-Specific Challenges
@@
| Benchmark | Task | Metric | System Prompt |
| Image Task | |||
| GQA [198] | CE-VQA | Exact Match | Answer the question using a single word or phrase. |
| MMB [199] | MC-VQA | Accuracy | Answer with the option’s letter from the given choices directly. |
| MME [200] | CE-VQA | Perception Score | Answer the question using a single word or phrase. |
| POPE [201] | CE-VQA | F1 Score | Answer the question using a single word or phrase. |
| ScienceQA-Image [202] | Visual reasoning | Exact Match | Answer with the option’s letter from the given choices directly. |
| SeedBench-Image [203] | MC-VQA | Accuracy | Answer with the option’s letter from the given choices directly. |
| VizWiz [204] | CE-VQA | Exact Match | When the provided information is insufficient, respond with “Unanswerable”. Answer the question using a single word or phrase. |
| VQA2 [205] | CE-VQA | Exact Match | Answer the question using a single word or phrase. |
| MM-Vet [206] | Visual reasoning | GPT-score | First please perform reasoning, and think step by step to provide the best answer to the following question: |
| LLaVA [1] | Visual reasoning | GPT-score | A chat between a curious human and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the human’s questions. |
| Video Task | |||
| ActivityNet [207] | CE-VQA | Accuracy / GPT-score | Answer the question using a single word or phrase. |
| VideoChatGPT [208] | OE-VQA | GPT-score | Evaluate the temporal accuracy of the prediction compared to the answer.* |
| NextQA [209] | CE-VQA | WUPS | Answer a question using a short phrase or sentence. |
| EgoSchema [210] | MC-VQA | Accuracy | Answer with the option’s letter from the given choices directly. |
| MVBench [8] | MC-VQA | Accuracy | Carefully watch the video and pay attention to the cause and sequence of events, the detail and movement of objects, and the action and pose of persons. Based on your observations, select the best option that accurately addresses the question. |
| LongVideo Bench [211] | MC-VQA | Accuracy | Answer with the option’s letter from the given choices directly. |
| VideoMME [212] | MC-VQA | Accuracy | Select the best answer to the following multiple-choice question based on the video and the subtitles. Respond with only the letter (A, B, C, or D) of the correct option. |
| PerceptionTest [213] | MC-VQA | Accuracy | Answer with the option’s letter from the given choices directly. |
| VideoDC [214] | Video Caption | GPT-score | Please provide a detailed description of the video, focusing on the main subjects, their actions, and the background scenes |
| AuroraCap [45] | Video Caption | VDCscore | Describe the video in detail. |
| Chardes-STA [215] | Temporal Grounding | IoU | Please find the visual event described by a sentence in the video, determining its starting and ending times. |
6.4.3 Deployment Hurdles
Despite their potential, many token compression methods face significant barriers to real-world deployment, stemming from a fundamental incompatibility with current large-scale model architectures and applications.
@@ 6.4.3 Deployment Hurdles
Furthermore, the design of certain techniques creates a fundamental disconnect with multi-turn conversational tasks. Methods that perform token compression internally within the model’s backbone or rely on cross-modal fusion are not natively compatible with this type of application. They lack an efficient mechanism to carry over and update a compressed representation across turns, instead requiring a costly re-computation of the entire conversation history for each new query.
6.4.4 Evaluation Challenges
@@
For methods requiring training, the interplay of various factors like training data and methodologies makes it challenging to isolate and directly compare the effectiveness of different approaches.
@@
Evaluation Benchmarks Gap. Current evaluation datasets for MLLM token compression, often relying on general multimodal benchmarks (Table IV), provide insufficient granularity. For example, in challenging long video understanding tasks, performance hinges more on sparse frame sampling capturing key frames than on the specific token compression method. This can obscure the true impact of token compression, making its efficacy appear negligible.
This reveals a critical gap: current datasets often fail to isolate and precisely measure the effect of token compression where it genuinely matters. Therefore, adopting evaluation methodologies similar to EffiVLM-Bench [216], which focuses on training-free acceleration evaluation, is crucial for accurately assessing the true efficacy and nuanced benefits of token compression methods.
6.5 Pruning Location and Trade-offs
@@
Pruning tokens at an early stage, such as within the encoder or projector, can dramatically shorten the sequence length. This significantly reduces the computational burden on the downstream LLM, leading to faster inference. However, this early compression carries a higher risk of discarding critical information, which can negatively impact overall model performance.
@@
6.6 Future Directions
6.6.1 Unified Multimodal Token Compression
Although different modalities exhibit unique redundancy patterns requiring specialized compression strategies, contemporary MLLMs predominantly focus on joint multimodal inference [10], whereas unimodal scenarios remain constrained by single-modality inputs. As established in Sections 3, 4, and 5, fundamental algorithmic principles (including transformation-based, similarity-based, attention-based, and query-based approaches) demonstrate transmodal applicability, indicating the viability of developing a unified multimodal token compression framework. Furthermore, leveraging cross-modal correlations represents a promising research direction for future token compression.
6.6.2 Improved Architecture
@@
7 Applications
The potential of multimodal token compression extends beyond technical enhancements, emerging as a universal efficiency engine for data-intensive AI systems. Multimodal models frequently process extreme-length token sequences exhibiting task-agnostic redundancy according to empirical analyses. Capitalizing on recent breakthroughs, we delineate four high-impact application domains:
7.1 GUI Agents and Human-Computer Interaction
Graphical user interface (GUI) agents perceive and interact with visual interfaces, interpret natural language instructions, analyze GUI states, and execute corresponding actions. These agents have to parse screen streams in real-time, producing extensive token sequences that often exceed computational limits [217, 148]. Multimodal token compression significantly enhances the efficiency of GUI agents. This approach mitigates context overflow in extended operation sequences by dynamically compressing redundant visual elements (e.g., extra white space or simple backgrounds). For some small but important control elements, it should also eliminate other irrelevant visual elements and highlight their importance. For instance, ShowUI [41] is the first model to apply token selection strategy to GUI agents. ShowUI segments GUI screenshots into connected components by clustering pixels with similar RGB values, significantly reducing the total number of discrete elements. During both training and inference phases, the system employs an adaptive token selection strategy that probabilistically prunes redundant tokens within these components, thereby optimizing computational efficiency while preserving functional semantics However, excessive compression risks inducing operational ambiguity, necessitating careful calibration.
7.2 Healthcare and Medical Imaging
Contemporary medical diagnosis and research critically depend on the interpretation and synthesis of multimodal medical imagery. MLLMs can integrate radiographic findings, medical histories, and ancillary diagnostic tests to generate differential diagnoses, which clinicians can correlate with patient records and physician notes to enhance diagnostic accuracy [218]. Furthermore, MLLMs can automatically draft preliminary radiology reports, potentially reducing the workload of radiologists [219, 220, 221]. However, high-dimensional modalities like PET-MRI exhibit prohibitively large data volumes while containing diagnostically critical yet sparse features [222]. Token compression techniques enable pathology-aware local lossless compression. Through hierarchical compression of anatomical structure tokens and pathological feature representations, these methods significantly reduce computational costs while maintaining diagnostic confidence levels. Critical attention must be devoted to preserving sub-resolution pathological features, particularly micro-lesions that may indicate early-stage conditions.
7.3 Robotics and Autonomous Systems

### v4

@@ IV-D1 Token Distillation
Token Turing Machines (TTMs) [104] maintain a compact external memory of summary tokens, sequentially compressing both new input tokens and memory at each timestep via a Transformer-based read/write mechanism, allowing scalable and efficient processing of long video sequences. BLIP-3-Video [105] introduces an explicit temporal encoder that abstracts hundreds of frame-level visual tokens into as few as 16–32 spatiotemporal tokens using learnable pooling and sequential models, enabling efficient video understanding with limited token usage. LinVT [108] proposes a plug-and-play Linear Video Tokenizer, which linearly aggregates frame-level visual tokens into a compact set of video tokens through spatio-temporal scoring, multi-scale pooling, and text-conditioned aggregation, enabling existing image-LLMs to efficiently process videos and dynamically extract question-relevant information. Long-VMNet [106] accelerates long-form video understanding by using a neural sampler to select discriminative visual tokens from clips and storing them in a fixed-size memory bank for each video; downstream queries are answered by processing only these memory tokens, greatly reducing computational cost while preserving key spatiotemporal information. STORM [107] inserts a Mamba-based [167] temporal encoder between the image encoder and LLM, using spatiotemporal scanning and pooling to inject temporal context into frame tokens and then aggressively compresses tokens by temporal and spatial pooling, enabling efficient long video understanding with minimal token loss. To understand more methods and applications of token distillation in video LLMs, please also refer to Section III-D for a detailed explanation.
IV-D2 Cross-Modal Selection
In video large language models (video LLMs), a query is commonly used to guide the selection of salient frames. In extreme cases, only a handful of frames are relevant to the posed question, allowing the tokens from the vast majority of remaining frames to be discarded. When dealing with an immense number of frames, finding query-relevant information can be akin to searching for a ”needle in a haystack” for the LLM. Query-based token compression methods can pre-filter query-relevant tokens, significantly alleviating the computational burden on the LLM.
@@ IV-D2 Cross-Modal Selection
V Audio-centric Token Compression
@@
The second category inherently reduces the number of tokens by the design of the tokenizer structure and the codebook. Nevertheless, detailed exploration of these specific design considerations falls outside the purview of this survey.
@@
V-A Transformation-based audio-centric Compression
@@
V-A1 Token Stacking
Similar to the pixel unshuffle operation in 2D image processing, this approach for audio LLMs token compression involves stacking multiple consecutive tokens along the hidden dimension of the token. This effectively reduces the total number of tokens. Notably, HTS-AT [109], an early example of audio token stacking for classification tasks within audio transformers, utilized 2D pixel-unshuffling on the 2D features extracted from Mel spectrograms to reduce audio tokens. More recent methods such as SLAM-ASR [110], LLaMA-Omni [111], Llama-AVSR [115] and others [172] stack the audio token. Since these token stacking operations alter the hidden dimension, an MLP is typically used to realign the dimension for compatibility with other modalities.
V-A2 Pooling
@@ V-A3 Temporal Convolution
V-B Similarity-based audio-centric Compression
Similarity-based compression methods aim for each audio token to carry unique information rather than being overly redundant. Similar to the ToMe [21] method used in vision transformers (ViT), A-ToMe [119] inserts a token merge module between the multihead self-attention (MHSA) and feed-forward network (FFN). This module merges adjacent audio tokens that have high cosine similarity.
V-C Attention-based audio-centric Compression
@@
V-C1 Attention in Encoder
Top-K [120] is a token selection method that operates within the audio spectrogram transformer block. It retains only the top K audio tokens, ranked by the magnitude of their attention scores. This prunes less attentive tokens, focusing on those with higher relevance as determined by the self-attention mechanism within the transformer.
V-C2 Attention in Decoder
@@
V-D Query-based audio-centric Compression
@@ V-D1 Token Distillation
Video-LLaMA [7] and SALMONN series [122, 123] employ an audio Q-former to transform variable-length audio inputs into a fixed-length sequence of learnable queries, thereby condensing audio information for the LLM. MMCE-Qformer [125] compresses acoustic information by utilizing learnable queries to extract global acoustic context from contextual audio embeddings. Concurrently, a cross-attention mechanism, guided by input text embeddings, captures local acoustic context relevant to each text token. This dual approach distills both broad and specific audio features into compact, text-relevant representations. MMS-LLaVA [126] reduces multimodal token length for efficient speech LLMs. It first halves the sequence length with an Early AV-Fusion Module, which combines visual and audio features. Subsequently, an AV Q-Former further compresses these fused features into a fixed number of queries, effectively capturing full speech context to bridge the token gap with text.
V-D2 Cross-Modal Selection
@@
VI Discussions
VI-A Synergies and Distinctions with Other Compression Methods
@@
VI-A1 Weight-Focused Compression Methods
These methods mainly target the model dimension () by reducing the effective size or complexity of the model weights. Model Quantization reduces weight precision, directly impacting the memory associated with . A key limitation is that highly aggressive quantization (e.g., 4-bit) often compromises accuracy, meaning there’s no ”free lunch” when it comes to achieving lossless performance. Furthermore, effectively accelerating these lower bit-rates often necessitates specialized hardware. Network Pruning removes redundant connections, effectively reducing the active parameters contributing to . For LLMs, aggressive structured pruning (e.g., beyond for downstream tasks) often leads to significant performance degradation or near-collapse due to the difficulty in preserving architectural integrity. Knowledge Distillation trains a smaller student model (with a smaller ) to mimic a larger teacher [183]. Its main limitation is the ”knowledge gap”, as the student may struggle to fully capture the teacher’s comprehensive knowledge, leading to performance disparities, especially on complex or out-of-distribution data. Low-Rank Factorization decomposes weight matrices into lower-rank approximations, thus reducing parameters related to . The challenge lies in finding an optimal low-rank approximation for diverse tasks without performance loss, as this is often task-dependent and complex to apply consistently across deep networks.
VI-A2 Token Compression
@@
VI-A3 Complementary Nature and Synergistic Gains
@@
VI-B Token Compression: Efficiency and Beyond
@@
VI-B1 Enhanced Modality Alignment
@@
VI-B2 Improved Information Representation
@@
Addressing these issues, the method classifications discussed earlier directly correspond to these types of data redundancy. Specifically, the transformation-based methods along with similarity-based approaches, are effective in mitigating the feature redundancy. Furthermore, attention-based methods play a crucial role in minimizing attention computation redundancy. Lastly, query-based methods are designed to reduce task-irrelevant redundancy.
VI-B3 Enable One-Shot Long-Context Understanding
@@
VI-C Combining Different Token Compression Methods
In Section VI-B2, we explored three distinct types of redundancy and the corresponding methods to reduce them. Since prune different forms of redundancy and can be applied at various stages, they seem orthogonal in their function. This raises a natural question: can we combine multiple token compression methods to achieve a synergistic effect?
Surprisingly, the answer is often no. Research has shown that a simple combination can lead to a phenomenon where the total effect is less than the sum of its parts. For example, the MoB [196] found that the combining methods do not exhibit superiority. This is probably attributed to the difference in each compression technique. When methods target different redundancies, their compression metrics can be inconsistent, leading to conflicting pruning decisions and a suboptimal outcome. From an information theory perspective, the amount of redundant information itself is finite, and different compression methods cannot repeatedly take effect on it. Therefore, while each method excels separately, their combined use does not guarantee a better result and may even degrade performance.
VI-D Current Challenges
VI-D1 Performance Degradation
While token compression can effectively condense multimodal features, it also introduces a risk of performance degradation. Current research on visual MLLMs, for example, shows that for models like LLaVA-OV-7B [2], near-lossless performance can be achieved by retaining as few as of the original tokens. However, performance declines sharply when the compression rate is pushed further. This challenge is more pronounced for larger and more recent models such as Qwen2.5-VL [33], LLaVA-Video-7B [34], and LLaVA-OV-72B [2], where achieving lossless compression seems to be more difficult.
This increased difficulty may stem from the models’ enhanced representational capabilities. It has been suggested that less capable models are inherently less sensitive to information loss from aggressive compression, as their weaker understanding already struggles to process the complex, uncompressed data fully. In contrast, more sophisticated models, which possess a more nuanced and holistic comprehension of multimodal tokens, are more susceptible to the subtle degradation caused by compression. For these models, achieving high performance requires a far more delicate and precise approach to preserve the token.
VI-D2 Task-Specific Challenges
@@
VI-D3 Deployment Hurdles
Despite their potential, many token compression methods face barriers to real-world deployment, stemming from a fundamental incompatibility with current large-scale model architectures and applications.
@@ VI-D3 Deployment Hurdles
Furthermore, task-aware token compression methods are not suit for multi-turn conversational tasks. Methods that perform token compression internally within the model’s backbone or rely on cross-modal fusion are not natively compatible with this type of application. They lack an efficient mechanism to carry over and update a compressed representation across turns, instead requiring a costly re-computation of the entire conversation history for each new query.
VI-D4 Evaluation Challenges
@@
For methods requiring training, various factors like training data and methodologies make it challenging to isolate and directly compare the effectiveness of different methods.
@@
Evaluation Benchmarks Gap. Current evaluation datasets for MLLM token compression, often relying on general multimodal benchmarks, provide insufficient granularity. For example, in challenging long video understanding tasks, performance hinges more on sparse frame sampling, capturing key frames, than on the specific token compression method. This can obscure the true impact of token compression, making its efficacy appear negligible.
This reveals a critical gap: current datasets often fail to isolate and precisely measure the effect of token compression. Therefore, adopting evaluation methodologies like EffiVLM-Bench [198], which focuses on training-free acceleration evaluation, is crucial for accurately assessing the true efficacy and nuanced benefits of token compression methods.
VI-E Pruning Location and Trade-offs
@@
Pruning tokens at an early stage, such as within the encoder or projector, can dramatically shorten the sequence length. This significantly reduces the computational burden on the downstream LLM, leading to faster inference. However, this early compression carries a higher risk of discarding critical information, which can negatively impact model performance.
@@
VI-F Future Directions
VI-F1 Unified Multimodal Token Compression
Although different modalities exhibit unique redundancy patterns requiring specialized compression strategies, contemporary MLLMs predominantly focus on joint multimodal inference [10], whereas unimodal scenarios remain constrained by single-modality inputs. As established in Sections III, IV, and V, fundamental algorithmic principles (including transformation-based, similarity-based, attention-based, and query-based approaches) demonstrate transmodal applicability, indicating the viability of developing a unified multimodal token compression framework. Furthermore, leveraging cross-modal correlations represents a promising research direction for the future.
VI-F2 Improved Architecture
@@
VII Applications
The potential of multimodal token compression extends beyond technical enhancements, emerging as a universal efficiency engine for data-intensive AI systems. Multimodal models frequently process extreme-length token sequences exhibiting high task-agnostic redundancy according to empirical analyses. Capitalizing on recent breakthroughs, we delineate four high-impact application domains:
VII-A GUI Agents and Human-Computer Interaction
Graphical user interface (GUI) agents perceive and interact with visual interfaces, interpret natural language instructions, analyze GUI states, and execute corresponding actions. These agents have to parse screen streams in real-time, producing extensive token sequences that often exceed computational limits [199, 148]. Multimodal token compression enhances the efficiency of GUI agents. This approach mitigates context overflow in extended operation sequences by dynamically compressing redundant visual elements (e.g., extra white space or simple backgrounds). For some small but important control elements, it should also eliminate other irrelevant visual elements and highlight their importance. For instance, ShowUI [41] is the first model to apply token selection strategy to GUI agents. ShowUI segments GUI screenshots into connected components by clustering pixels with similar RGB values, significantly reducing the total number of discrete elements. During both training and inference phases, the system employs an adaptive token selection strategy that probabilistically prunes redundant tokens within these components, thereby optimizing computational efficiency while preserving functional semantics However, excessive compression risks inducing operational ambiguity, necessitating careful calibration.
VII-B Healthcare and Medical Imaging
The effective synthesis of multimodal medical data is pivotal to advancing contemporary medical diagnosis and research. MLLMs can integrate radiographic findings, medical histories, and ancillary diagnostic tests to generate differential diagnoses, which clinicians can correlate with patient records and physician notes to enhance diagnostic accuracy [200]. Furthermore, MLLMs can automatically draft preliminary radiology reports, potentially reducing the workload of radiologists [201, 202, 203]. A major challenge for MLLMs in medical imaging is the processing of high-resolution images, such as Whole-slide Images (WSIs) in pathology, which can contain billions of pixels. To overcome this, TCP-LLaVA [204] uses a set of trainable compression tokens to aggregate and condense crucial information from thousands of visual and textual inputs. Instead of feeding every single image patch token into the language model, only these compressed tokens are forwarded for answer generation. Token compression holds vast potential for widespread application in the field of healthcare and medical imaging, enabling the efficient analysis of complex, high-resolution images.
VII-C Robotics and Autonomous Systems